Grade 10Mechanics

in a hydraulic lift the surface area of the input piston is 10cm2 the surface area of the output piston is 3000 cm2 A 100N forcee applied to the input piston raises the output piston by1m calculate the work done .

Aiden Mathew Denny
8 Years agoGrade 10

1 Answer

Arun
8 Years ago
Pressure input on piston
P = F/A = 100 /10* 10^-4
 
P = 10^5 N/m²
 
According to Pascal's law
 
10^5 = fo /3000* 10^-4
 
fo = 3 *10^4
 
Now
Work done = 3 *10^4 * 1 = 3 *10^4 joules